Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

GMAT Pre Questions ---Probability

Expert replies
by maypoon » Sun Feb 08, 2009 8:47 am
Dear all,

Please help me to deal with the following questions related to "Probabilty".
Thank you so much!


Q1. From a group of 3 boys & 4 girls, 4 children are to be randomly selected. What's the probabilty that equal numbers of boys & grils will be selected?

A. 1/10
B. 4/9
C. 1/2
D. 3/5
E. 2/3

Ans D


Q.2. There're 11 women & 9 men in a certain club. If the club is to select a committe of 2 women & 2 men, how many different such committees are possible?

A. 120
B. 720
C. 1060
D.1250
E.1980

Q3. A shipment of 8 television sets contains 2 black -and white sets & 6 colors sets . If 2 television sets are to be chosen at random from this shipment , what 's the probabilty that at least 1 of the 2 sets chosen will be black -and -white sets?

A. 1/7
B. 1/4
C.5/14
D.11/28
E.13/26

Ans. E


Best Wishes,
May
Join the discussion
Source: — Problem Solving |

by earth@work » Sun Feb 08, 2009 12:04 pm
Q1.
Soln: Choosing 2 boys out of 3 boys=3C2=3!/2! = 3 ways
Choosing 2 girls out of 4 girls = 4C2=6 ways
Total Favourable ways = 3*6=18
Now, total ways = 7C4=7!/4!*3! = 35
P(2B& 2G) = 18/35 ..... my answer does not match ans D

Q.2.
Soln: Choosing 2 women out of 11 = 11C2 = 55
choosing 2 men out of 9 =9C2 = 36
total no . committees = 55*36 = 1980 - Ans E

Q3.
Prob(atleast 1 B/W tv) = 1- Prob(NO b/w tv)
Prob of no b/w tv = prob of both Colour tv
P(colour tv) = 6C2/8C2 =15/28
P(atleast one b/w tv) = 1-15/28 = 13/28 (this is again not matching your answer choice E 13/26 )
COuld u pls confirm the answers for 1 & 3
Join the discussion

by betamax » Sun Feb 08, 2009 12:22 pm
I got the same.
1. 18/35
2. 1980
3. 13/28
I would like to see the solutions to these.
Join the discussion

Dear all,

I also have no idea about the ans of Q1 & Q3.

The ans of mine of Q3 is also 13/28.

And for the Q1, I have totally no idea !!

Anyway, thank you so much for your help!

If I find out the solution..I will post it later !!

Best Wishes,
May :D
Join the discussion

by BuckeyeT » Mon Feb 09, 2009 8:17 am
May,

I'm getting the same 18/35, albeit by a slightly different method.

Q1: Probability of equal boys and girls on 4 picks. First, how many possible selection choices are there? Write out each possible choice possibility that meets the requirement:
C1 = B,B,G,G
C2 = B,G,B,G
C3 = B,G,G,B
C4 = G,B,G,B
C5 = G,B,B,G
C6 = G,G,B,B

So to determine the probability of these choices, we need to determine the probability of any choice type and multiply by 6.

C1 = B,B,G,G
P(C1) = P(1st pick boy)*P(2nd pick boy)*P(3rd pick girl)*P(4th pick girl)
P(C1) = (3/7)*(2/6)*(4/5)*(3/4) = 3/35

P(C1) * 6 = (3/35)*6 = 18/35
Join the discussion

by ashishsj » Wed Feb 25, 2009 8:35 am
The actual question is:

From a group of 3 boys & 3 girls, 4 children are to be randomly selected. What's the probabilty that equal numbers of boys & grils will be selected?

Ans: 3/5
Join the discussion